# Option Pricing with Time-Varying Volatility Risk Aversion

### v3

@@ 2.1 Pricing Kernel with Constant Parameters (CHNG)
$\frac{M_{t+1}}{M_{t}}=\left(\frac{S_{t+1}}{S_{t}}\right)^{\phi}\exp\left[\delta+\pi h_{t+1}+\xi(h_{t+2}-h_{t+1})\right],$ (3)
@@
$M_{t+1,t}\equiv\frac{M_{t+1}}{\mathbb{E}_{t}\left[M_{t+1}\right]}=\frac{\exp\left(\phi R_{t+1}+\xi h_{t+2}\right)}{\mathbb{E}_{t}[\exp(\phi R_{t+1}+\xi h_{t+2})]},$ (4)
where the conditional expectation in the denominator is taken with respect to the natural filtration for returns under $\mathbb{P}$, i.e. $\mathbb{E}_{t}(\cdot)=\mathbb{E}^{\mathbb{P}}(\cdot|R_{t},R_{t-1},\ldots)$. The expression above shows that the pricing kernel depends on both equity risk and variance risk, where the latter is typically characterized by the parameter $\xi$. The equity risk is governed by $\phi$, as well as $\xi$, because $h_{t+2}$ depends on return $R_{t+1}$. This structure implies the following risk-neutral dynamics:
@@
with $z_{t+1}^{*}\sim iid\ N(0,1)$ and the following relationships between parameters under $\mathbb{P}$ and $\mathbb{Q}$
$h_{t}^{*}=h_{t}\eta,\quad\omega^{*}=\omega\eta,\quad\beta^{\ast}=\beta,\quad\alpha^{*}=\alpha\eta^{2},\quad\gamma^{\ast}=\tfrac{1}{\eta}(\gamma+\lambda-\tfrac{1}{2})+\tfrac{1}{2},\quad\eta\equiv(1-2\alpha\xi)^{-1}.$ (5)
@@
where $\kappa_{0}=-\frac{1}{2}\log\eta$ and $\kappa_{1}=\frac{1}{2}(\lambda-\tfrac{1}{2})^{2}-\frac{1}{2}\eta$ [^2] The expressions for $\kappa_{0}$ and $\kappa_{1}$ in Christoffersen et al., (2013) are $\kappa_{0}=\delta+\xi\omega+\phi r$ and $\kappa_{1}=\pi+\xi(\beta-1+\alpha(\lambda-\tfrac{1}{2}+\gamma)^{2})$. It follows from our results in Lemma 1 that they are equivalent to those presented here. Expression (6) shows that $\lambda$ and $\xi$ define the linear and quadratic terms, respectively, with $\xi>0$ leading to a U-shaped and $\xi<0$ to an inverted U-shaped pricing kernel.
Empirically, this relationship tends to be U-shaped, but it is unstable over time as can be seen in Christoffersen et al., (2013, figure 3). Figure 1 includes parts of figures 3 and 5 in Christoffersen et al., (2013); the four left panels are based on model-free market prices and the corresponding four right panels are based on model prices using their estimated CHNG model.
@@
(b) Model-based log ratios
Figure 1: Log ratios of risk-neutral one-month densities and physical one-month histogram against log returns in monthly standard deviations. Panels (a) and (b) are subsets of figures 3 and 5, respectively, in Christoffersen et al., (2013). The former is based on empirical market prices and the latter has the corresponding model-based log ratios.
@@ 2.2 Dynamic Pricing Kernel by Means of Variance Risk Aversion
We seek a more flexible pricing kernel that is coherent with the observed variation over time. Next, we introduce a new dynamic pricing kernel while maintaining the Heston-Nandi GARCH structure under physical measure, $\mathbb{P}$.
@@
where $\phi_{t}$ and $\xi_{t}$ are $\mathcal{G}_{t}$-measurable, and $\mathbb{E}_{t}^{\mathbb{P}}(\cdot)\equiv\mathbb{E}^{\mathbb{P}}(\cdot|\mathcal{G}_{t})$.
@@
$(ii)$ The return dynamics under the risk-neutral measure $\mathbb{Q}$ is given by:
$R_{t+1}$ $=$ $r-\tfrac{1}{2}h_{t+1}^{*}+\sqrt{h_{t+1}^{*}}z_{t+1}^{*},$ (8)
@@
with $z_{t+1}^{*}\sim iid\ N(0,1)$ and the following relations between $\mathbb{P}$ and $\mathbb{Q}$ parameters:
@@
$(iii)$ The parameter, $\eta_{t}$, equals the ratio of the conditional variances under $\mathbb{Q}$ and $\mathbb{P}$,
$\eta_{t}=\frac{h_{t+1}^{\ast}}{h_{t+1}}.$ (9)
@@
First, $\eta_{t}=h_{t+1}^{\ast}/h_{t+1}$ emerges as a fundamental quantity, which is termed the variance risk ratio. [^3] Parameterize the model with the inverse ratio, $\theta_{t}=h_{t+1}/h_{t+1}^{\ast}$, simplifies several expressions, such as $\xi_{t}=\tfrac{1}{2\alpha}(1-\theta_{t})$ and $\phi_{t}=(1-\theta_{t})(\gamma-\tfrac{1}{2})-\theta_{t}\lambda$. We adopt $\eta_{t}$, because it is more common in the literature. This variable appears in all expressions relating $\mathbb{P}$ and $\mathbb{Q}$, and it has a straightforward interpretation. Its functional relationship with $\xi_{t}$ shows that $\eta_{t}$ is an indirect measure of volatility risk aversion. When $\eta_{t}$ is large, agents demand a large compensation for taking on variance risk, while a small value of $\eta_{t}$ corresponds to an appetite for variance risk. Moreover, $\eta_{t}$ is obviously related to the VRP, where the latter concerns the expectation of future values of $h_{t+1}-h_{t+1}^{\ast}$. This follows from $h_{t+1}-h_{t+1}^{\ast}=\left(1-\eta_{t}\right)h_{t+1}$. This relation could be used to construct a model-based measure of the VRP. Note that the VRP can be positive in the most general version of the model. A negative VRP can be guaranteed in the model design by restricting $\eta_{t}$ to be greater than one. [^4] This can be be achieved by specifying a dynamic model for $\log(\eta_{t}-1)$. We do not impose this restriction in our analysis, as we specify a model for $\log\eta_{t}$. For comparison, the ratio, $h_{t}^{\ast}/h_{t}$, is constant in CHNG, and the relation $h_{t}-h_{t}^{\ast}=(1-\eta)h_{t}$ shows that CHNG implicitly restricts the VRP to be proportional to the conditional variance under $\mathbb{P}$.
@@
$\log M_{t+1,t}$ $=-\lambda R_{t+1}+\xi_{t}\alpha\frac{\left(R_{t+1}-r\right)^{2}}{h_{t+1}}+\kappa_{0,t}+\kappa_{1,t}h_{t+1},$ (10)
is equivalent to (7), where $\kappa_{0,t}=-\frac{1}{2}\log\eta_{t}$, and $\kappa_{1,t}=\tfrac{1}{2}\left(\lambda-\tfrac{1}{2}\right)^{2}-\tfrac{1}{2}\eta_{t}$.
@@
Fourth, an important reason for permitting $\eta_{t}$ to be time-varying is that it enables $h_{t}^{\ast}$ to incorporate information from derivative prices. In conventional GARCH models, the conditional variance depends only on lagged returns, such that $h_{t+1}\in\sigma(R_{t},R_{t-1},\ldots)$. A restrictive implication of a constant VRR is that it requires the two conditional variances to be proportional to each other. Consequently, $h_{t+1}^{\ast}$ is driven solely by lagged returns, leaving no room for derivative prices to influence its dynamic properties. Allowing $\eta$ to be time-varying eliminates this rigidity and enables the model to incorporate information from derivative prices into the dynamic model of $h_{t+1}^{\ast}$. We take advantage of this feature in our empirical implementation.
@@
Sixth, the structure introduced in Theorem 1 generates time-varying leverage under $\mathbb{Q}$, which manifests itself in a number of ways. A time-varying leverage effect will impact the distribution of cumulative returns, and the lower panels of Figure 2 show the skewness and kurtosis of cumulative returns under $\mathbb{Q}$ as a function of the number of days that returns are cumulated over (the $x$-axis). Skewness and kurtosis are shown for the two levels of the VRR, $\eta_{\mathrm{low}}$ and $\eta_{\mathrm{high}}$, and these result in distinctively different levels of skewness and kurtosis under $\mathbb{Q}$. Both skewness and kurtosis are far more pronounced for the large value of $\eta$. Another way to illustrate the leverage effect is with the news impact curve by Engle and Ng, (1993). The news impact curve under $\mathbb{Q}$ is defined by $\mathrm{NIC}(z^{\ast})=\mathbb{E}_{t}^{\mathbb{Q}}(h_{t+1}^{\ast}|z_{t}^{\ast}=z^{\ast})-\mathbb{E}_{t}^{\mathbb{Q}}(h_{t+1}^{\ast}|z_{t}^{\ast}=0)$, and from Theorem 1 it follows that
$\mathrm{NIC}(z^{\ast})=\alpha\eta_{t}\eta_{t-1}z^{\ast 2}-2\alpha\gamma\eta_{t}\sqrt{h_{t}^{*}}z^{\ast},$ (11)
whose shape depends on the VRR. The upper right panel of Figure 2 presents the news impact curve, with $h_{t}^{\ast}$ equal to its unconditional mean under $\mathbb{Q}$, as estimated in Section 5, and with $\eta=\eta_{t}=\eta_{t-1}$ equal to either $\eta_{\mathrm{low}}$ (red dashed line) or $\eta_{\mathrm{high}}$ (solid blue line). We observe the asymmetric shape of the news impact curve, see e.g. Chen and Ghysels, (2011). However, the level of asymmetry depends on the level of $\eta$, such that volatility is more responsive to negative return shocks when $\eta$ is large, than when $\eta$ is small.
@@ 3.1 VIX Pricing
${\rm VIX}_{t}=A\times\sqrt{\frac{1}{M}\sum_{k=1}^{M}\mathbb{E}_{t}^{\mathbb{Q}}(h_{t+k}^{*})},$
@@
Suppose that $\log\eta_{t}$ is an autoregressive moving average process of orders $(p,q)$,
@@
where $\varphi(z)=0\Rightarrow|z|>1$, $\theta(z)$ and $\varphi(z)$ have no common roots, $\{\varepsilon_{t}\}$ is a sequence of iid random variables for which, $\mathbb{E}^{\mathbb{Q}}(\varepsilon_{t}|\mathcal{G}_{t-1})=0$, $\mathrm{var}^{\mathbb{Q}}(\varepsilon_{t}|\mathcal{G}_{t-1})=\sigma^{2}$, and its MGF satisfies $\operatorname{mgf}_{\varepsilon}(s)\equiv\mathbb{E}^{\mathbb{Q}}\left[\exp\left(s\varepsilon_{t}\right)|\mathcal{G}_{t-1}\right]<\infty$ a.s. for $s\in\mathbb{R}$.
Assumption 3 is explicit about the first two conditional moments of $\varepsilon_{t}$, but does not impose restrictive distributional assumptions. However, the exact distribution is important for VIX and option prices, as these depend on $\operatorname{mgf}_{\varepsilon}(s)$.
@@
Suppose that Assumptions 1-3 hold, then the VIX is given by
@@
where $a_{1}(M,\sigma^{2})$ and $a_{2}(M,\sigma^{2})$ depend on $(\eta_{t},\ldots,\eta_{t-p+1})$, $(\varepsilon_{t},\ldots,\varepsilon_{t-q+1})$, and $\operatorname{mgf}_{\varepsilon}(s)$. See Appendix A.3 for their exact expressions. [^6] We have suppressed the terms dependence on $\varphi$, $\zeta$, and the parameters of the HNG model and (13) suppresses an approximation term that is negligible in practice (see Lemma A.1).

### v4

@@ 1.1 Pricing Kernel with Constant Parameters (CHNG)
$\frac{M_{t+1}}{M_{t}}=\left(\frac{S_{t+1}}{S_{t}}\right)^{\phi}\exp\left[\delta+\pi h_{t+1}+\xi(h_{t+2}-h_{t+1})\right],$
@@
$M_{t+1,t}\equiv\frac{M_{t+1}}{\mathbb{E}^{\mathbb{P}}[M_{t+1}|\mathcal{F}_{t}^{R}]}=\frac{\exp\left(\phi R_{t+1}+\xi h_{t+2}\right)}{\mathbb{E}^{\mathbb{P}}[\exp(\phi R_{t+1}+\xi h_{t+2})|\mathcal{F}_{t}^{R}]},$
where $\mathcal{F}_{t}^{R}=\sigma\left(\{R_{j}\},j\leq t\right)$ is the natural filtration for returns only. The pricing kernel above depends on both equity risk and variance risk, where the latter is typically characterized by the parameter $\xi$. The equity risk is governed by $\phi$, as well as $\xi$, because $h_{t+2}$ depends on return $R_{t+1}$. This pricing kernel implies the following risk-neutral dynamics:
@@
with $z_{t+1}^{*}|\mathcal{F}_{t}^{R}\overset{\mathbb{Q}}{\sim}iid\ N(0,1)$ and the following relations between $\mathbb{P}$ and $\mathbb{Q}$ parameters:
$h_{t}^{*}=h_{t}\eta,\quad\omega^{*}=\omega\eta,\quad\beta^{\ast}=\beta,\quad\alpha^{*}=\alpha\eta^{2},\quad\gamma^{\ast}=\tfrac{1}{\eta}(\gamma+\lambda-\tfrac{1}{2})+\tfrac{1}{2},\quad\eta\equiv(1-2\alpha\xi)^{-1}.$
@@
where $\kappa_{0}=-\frac{1}{2}\log\eta$ and $\kappa_{1}=\frac{1}{2}(\lambda-\tfrac{1}{2})^{2}-\tfrac{1}{8}\eta$ [^2] The expressions for $\kappa_{0}$ and $\kappa_{1}$ in Christoffersen et al., (2013) are $\kappa_{0}=\delta+\xi\omega+\phi r$ and $\kappa_{1}=\pi+\xi(\beta-1+\alpha(\lambda-\tfrac{1}{2}+\gamma)^{2})$. It follows from our results in Lemma 1 that they are equivalent to those presented here. Expression (3) shows that $\lambda$ and $\xi$ define the linear and quadratic terms, respectively, with positive $\xi$ leading to a U-shaped and negative $\xi$ to an inverted U-shaped pricing kernel. Empirically, this relationship tends to be generally U-shaped, but it is unstable over time as seen in Christoffersen et al., (2013, figure 3). Figure 1 includes parts of figures 3 and 5 in Christoffersen et al., (2013); the four left panels are based on model-free market prices and the corresponding four right panels are based on model prices using their estimated CHNG model.
@@
(b) CHNG model-based log ratios
Figure 1: Log ratios of risk-neutral one-month densities and physical one-month histogram against log returns in monthly standard deviations. Panels (a) and (b) are subsets of figures 3 and 5, respectively, in Christoffersen et al., (2013). The former is based on empirical market prices and the latter has the corresponding CHNG model-based log ratios.
Alt text: The shape of the pricing kernel based on market prices and the CHNG model for four calendar years. The figure plots log ratios of risk-neutral to physical one-month densities against log returns. The shapes for 2001 and 2002 are similar, while those for 2005 and 2006 are distinctively different.
@@ 1.2 Dynamic Pricing Kernel by Means of Variance Risk Aversion
We seek a more flexible pricing kernel that is coherent with the observed variation over time. Next, we introduce a new dynamic pricing kernel while maintaining the Heston-Nandi GARCH structure (1)-(2) under physical measure, $\mathbb{P}$.
@@
where $\phi_{t}$ and $\xi_{t}$ are $\mathcal{G}_{t}$-measurable.
@@
$(ii)$ The return dynamics under the risk-neutral measure $\mathbb{Q}$ are given by:
$R_{t+1}$ $=$ $r-\tfrac{1}{2}h_{t+1}^{*}+\sqrt{h_{t+1}^{*}}z_{t+1}^{*},$
@@
where $z_{t+1}^{*}|\mathcal{G}_{t}\overset{\mathbb{Q}}{\sim}iid\ N(0,1)$ and the following relations between $\mathbb{P}$ and $\mathbb{Q}$ parameters:
@@
$(iii)$ The dynamic parameter, $\eta_{t}$, equals the ratio of the conditional variances under $\mathbb{Q}$ and $\mathbb{P}$,
$\eta_{t}=\frac{h_{t+1}^{\ast}}{h_{t+1}}.$
@@
First, $\eta_{t}=h_{t+1}^{\ast}/h_{t+1}$ emerges as a fundamental quantity [^3] Parameterize the model with the inverse ratio, $\theta_{t}=h_{t+1}/h_{t+1}^{\ast}$, simplifies several expressions, such as $\xi_{t}=\tfrac{1}{2\alpha}(1-\theta_{t})$ and $\phi_{t}=(1-\theta_{t})(\gamma-\tfrac{1}{2})-\theta_{t}\lambda$. We adopt $\eta_{t}$, because it is more common in the literature., which is termed the variance risk ratio. It is $\mathcal{G}_{t}$-measurable since $\eta_{t}$ is a function of $\xi_{t}\in\mathcal{G}_{t}$ defined in Assumption 2. This variable appears in all expressions relating $\mathbb{P}$ and $\mathbb{Q}$, and it has a straightforward interpretation. Its functional relationship with $\xi_{t}$ shows that $\eta_{t}$ is an indirect measure of volatility risk aversion. When $\eta_{t}$ is large, agents demand a large compensation for taking on variance risk, while a small value of $\eta_{t}$ corresponds to an appetite for variance risk. Moreover, $\eta_{t}$ is obviously related to the VRP, where the latter concerns the expectation of future values of $h_{t+1}-h_{t+1}^{\ast}$. This follows from $h_{t+1}-h_{t+1}^{\ast}=\left(1-\eta_{t}\right)h_{t+1}$. This relation could be used to construct a model-based measure of the VRP. Note that the VRP can be positive in the most general version of the model. A negative VRP can be guaranteed in the model design by restricting $\eta_{t}$ to be greater than one. [^4] This can be be achieved by specifying a dynamic model for $\log(\eta_{t}-1)$. We do not impose this restriction in our analysis, as we specify a model for $\log\eta_{t}$. For comparison, the ratio, $h_{t}^{\ast}/h_{t}$, is constant in CHNG model, and the relation $h_{t}-h_{t}^{\ast}=(1-\eta)h_{t}$ shows that CHNG model implicitly restricts the VRP to be proportional to the conditional variance under $\mathbb{P}$.
@@
$\log M_{t+1,t}$ $=-\lambda R_{t+1}+\xi_{t}\alpha\frac{\left(R_{t+1}-r\right)^{2}}{h_{t+1}}+\kappa_{0,t}+\kappa_{1,t}h_{t+1},$
is equivalent to (4), where $\kappa_{0,t}=-\frac{1}{2}\log\eta_{t}$, and $\kappa_{1,t}=\tfrac{1}{2}(\lambda-\tfrac{1}{2})^{2}-\tfrac{1}{8}\eta_{t}$.
@@
Alt text: Properties of cumulative returns for two levels of the variance risk ratio: low and high. The figure consists of four panels displaying (1) the log-ratio of risk-neutral to physical densities, (2) news impact curves, (3) skewness under the risk-neutral measure, and (4) kurtosis under the risk-neutral measure.
@@
Fourth, an important reason for permitting $\eta_{t}$ to be time-varying is that it enables $h_{t}^{\ast}$ to incorporate information from derivative prices. In conventional GARCH models, the conditional variance depends only on lagged returns, such that $h_{t+1}\in\mathcal{F}_{t}^{R}$. A restrictive implication of a constant VRR is that it requires the two conditional variances to be proportional to each other. Consequently, $h_{t+1}^{\ast}$ is driven solely by lagged returns, leaving no room for derivative prices to influence its dynamic properties. Allowing $\eta$ to be time-varying eliminates this rigidity and enables the model to incorporate information from derivative prices into the dynamic model of $h_{t+1}^{\ast}$. We take advantage of this feature in our empirical implementation.
@@
Sixth, the structure introduced in Theorem 1 generates time-varying leverage under $\mathbb{Q}$, which manifests itself in a number of ways. A time-varying leverage effect will impact the distribution of cumulative returns, and the lower panels of Figure 2 show the skewness and kurtosis of cumulative returns under $\mathbb{Q}$ as a function of the number of days that returns are cumulated over (the $x$-axis). Skewness and kurtosis are shown for the two levels of the VRR, $\eta_{\mathrm{low}}$ and $\eta_{\mathrm{high}}$, and these result in distinctively different levels of skewness and kurtosis under $\mathbb{Q}$. Both skewness and kurtosis are far more pronounced for the large value of $\eta$. Another way to illustrate the leverage effect is with the news impact curve by Engle and Ng, (1993). The news impact curve under $\mathbb{Q}$ is defined by $\mathrm{NIC}(z^{\ast})=\mathbb{E}_{t}^{\mathbb{Q}}(h_{t+1}^{\ast}|z_{t}^{\ast}=z^{\ast})-\mathbb{E}_{t}^{\mathbb{Q}}(h_{t+1}^{\ast}|z_{t}^{\ast}=0)$, and from Theorem 1 it follows that $\mathrm{NIC}(z^{\ast})=\alpha\eta_{t}\eta_{t-1}z^{\ast 2}-2\alpha\gamma\eta_{t}\sqrt{h_{t}^{*}}z^{\ast}$, whose shape depends on the VRR. The upper right panel of Figure 2 presents the news impact curve, with $h_{t}^{\ast}$ equal to its unconditional mean under $\mathbb{Q}$, as estimated in Section 4, and with $\eta=\eta_{t}=\eta_{t-1}$ equal to either $\eta_{\mathrm{low}}$ (dashed line) or $\eta_{\mathrm{high}}$ (solid line). We observe the asymmetric shape of the news impact curve, see e.g. Chen and Ghysels, (2011). However, the level of asymmetry depends on the level of $\eta$, such that volatility is more responsive to negative return shocks when $\eta$ is large, than when $\eta$ is small.
@@ 2.1 VIX Pricing
${\rm VIX}_{t}=A\times\sqrt{\frac{1}{M}\sum_{k=1}^{M}\mathbb{E}_{t}^{\mathbb{Q}}(h_{t+k}^{*})},$ (5)
@@
Suppose that the logarithm of variance risk ratio, $\log\eta_{t}$, follows an autoregressive moving average process of orders $(p,q)$,
@@
where $\varphi(x)=0\Rightarrow|x|>1$, $\theta(x)$ and $\varphi(x)$ have no common roots. The innovation $\varepsilon_{t}$ is $\mathcal{G}_{t}$-measurable, and $\varepsilon_{t+1}|\mathcal{G}_{t}$ follows an iid distribution with $\mathbb{E}_{t}^{\mathbb{Q}}(\varepsilon_{t+1})=0$, $\mathrm{var}_{t}^{\mathbb{Q}}(\varepsilon_{t+1})=\sigma^{2}$, and a MGF satisfying $\operatorname{mgf}_{\varepsilon}(s)\equiv\mathbb{E}_{t}^{\mathbb{Q}}\left[\exp\left(s\varepsilon_{t+1}\right)\right]<\infty$ a.s. for $s\in\mathbb{R}$. Additionally, $\varepsilon_{t}$ is independent of the return shock $z_{\tau}$ for all $\tau$.
Assumption 3 is explicit about the first two conditional moments of $\varepsilon_{t}$, but does not impose restrictive distributional assumptions. However, the exact distribution is important for VIX and option prices, as these depend on $\operatorname{mgf}_{\varepsilon}(s)$. That $\varepsilon_{t}$ is $\mathcal{G}_{t}$-measurable follows from Assumption 2 and therefore not a new assumption.
@@
Suppose that Assumptions 1-3 hold, then the analytical expression for $M$-period ahead VIX in (5) is given by
@@
where $a_{1}(M,\sigma^{2})$ and $a_{2}(M,\sigma^{2})$ both depend on $(\eta_{t},\ldots,\eta_{t-p+1})$, $(\varepsilon_{t},\ldots,\varepsilon_{t-q+1})$, and $\operatorname{mgf}_{\varepsilon}(s)$. See Appendix A.3 for their exact expressions. [^6] We have suppressed the terms dependence on $\varphi$, $\zeta$, and the parameters of the HNG model and (7) suppresses an approximation term that is negligible in practice (see Lemma A.1).